OpenStudy (anonymous):

Is 36x^4-64x^2 a special case polynomial?

OpenStudy (mathstudent55):

Notice that 36x^4 is the square of 6x^2, and 64x^2 is the square of 8x, so you have a difference of two squares binomial.

OpenStudy (anonymous):

Thanks that is a big help

OpenStudy (anonymous):

Can you solve 36x^4-64x^2 for me? I'm confused

OpenStudy (mathstudent55):

Do you need to factor it?

OpenStudy (anonymous):

I think so? It just says to show my work.

OpenStudy (mathstudent55):

a^2 - b^2 = (a + b)(a - b) 36x^4 - 64x^2 First step is to try to factor a common factor: 4x^2(9x^2 - 16) = 4x^2(3x + 4)(3x - 4)

OpenStudy (anonymous):

thanks, will you help me with another one?

OpenStudy (mathstudent55):

sure. pls start a new post. i'll look for it

OpenStudy (anonymous):

where do I do that? I can't see where to do it:(